# ConceptDrift: Uncovering Biases through the Lens of Foundation Models

### v2

@@ Abstract
An important goal of ML research is to identify and mitigate unwanted biases intrinsic to datasets and already incorporated into pre-trained models. Previous approaches have identified biases using highly curated validation subsets, that require human knowledge to create in the first place. This limits the ability to automate the discovery of unknown biases in new datasets. We solve this by using interpretable vision-language models, combined with a filtration method using LLMs and known concept hierarchies. More exactly, for a dataset, we use pre-trained CLIP models that have an associated embedding for each class and see how it drifts through learning towards embeddings that disclose hidden biases. We call this approach ConceptDrift and show that it can be scaled to automatically identify biases in datasets like ImageNet without human prior knowledge. We propose two bias identification evaluation protocols to fill the gap in the previous work and show that our method significantly improves over SoTA methods, both using our protocol and classical evaluations. Alongside validating the identified biases, we also show that they can be leveraged to improve the performance of different methods. Our method is not bounded to a single modality, and we empirically validate it both on image (Waterbirds, CelebA, ImageNet), and text datasets (CivilComments).
@@ 1 Introduction
Deep neural networks, and especially fine-tuned versions of foundation models, are commonly deployed in critical areas such as healthcare, finance, and criminal justice, where biased predictions can have significant societal consequences [1]. Despite their impact, these models are often employed in their natural black-box state, i.e. as highly non-linear, multi-layered decision processes, lacking transparency or interpretability. Even if the pretrained model has been validated by the community, the dataset leveraged in the fine-tuning process can, and usually does, imprint the model with new biases. This issue is particularly concerning as biases can lead to undesired outcomes [6], reinforcing existing inequalities or creating new forms of discrimination.
Subpopulation shift setups [40, 21, 5], featuring naturally occurring spurious correlations, provided researchers with a controlled environment for studying bias detection and prevention. Within this context, recent efforts began employing foundation models [15, 47] in their investigation protocols. However, current methods limit themselves to data analysis alone. For instance, the method proposed by Kim et al. [15] relies on validating the presence of a given object within the set of mistakes and its absence from within the set of the correctly classified samples within the validation set, in order to label it as a bias. Thus, data-restricted methods rely on the critical assumption that the validation set is exhaustive in some sense, that it contains samples capable of exposing all biases. This assumption is justified only within the context of a controlled experimental environment, such as having a subpopulation shift setup where all sources of bias are given or known. Any other real-world scenario hardly guarantees an exhaustive validation set.
@@
Figure 1: Following ConceptDrift’s steps for the ImageNet ”Fire Truck” class. In Step 1, during training, the classification weights $W$ drift from the initial class concept embedding $W^{0}$, outside the scope of relevant concepts, towards biases. In Step 2, our method filters out class-related concepts and, using an embedding-space scoring system, ranks and automatically selects the highest-ranking class-neutral concepts as biases. Leveraging these identified biases, we generate synthetic images that are able to fool even strong pre-trained methods.
@@
We expand the current usage of foundation models beyond the restricted scope of data analysis and propose a new direction for bias identification. Our method focuses on investigating the skewness of the model’s weights towards detecting and prioritizing spurious features as part of the decision-making process of the investigated model. We hereby propose a novel protocol for uncovering biases using foundation models such as CLIP [30] and mGTE [45], leveraging the topology of their embedding space to identify and name biases instilled by linear probing. Our protocol, dubbed ConceptDrift, is illustrated in Fig. 1. We showcase how, during training, the weights of the final classification layer drift away from the textual representation of their associated class, towards representations of spurious attributes. We propose a ranking system based on embedding-space arithmetic to extract concepts that factor in the activation of class neurons and delineate the highest-ranking concepts that lie outside the semantic scope of the classes, as biases.
Our main contributions are summarized as follows:
@@
    We introduce ConceptDrift, a method capable of pin-pointing concepts that are relevant for the decision-making process of a model. We are the first to propose a weight-space approach for identifying biases that a foundation model learns from a given dataset, diverging from the current techniques oriented on validation set errors.
@@
    We propose two novel evaluation strategies for the bias identification task. The first one evaluates the performance decay over generated samples that exploit the identified biases, highlighting the potential harm biases can cause to models. The second one assesses the performance gains achieved through bias-based regularization in a training setup with a 100% spurious correlation between classes and biases, focusing on mitigating biases and enhancing model robustness.
@@
    We validate ConceptDrift, using both established evaluation protocols and our newly introduced strategies. We achieve significant advances upon the SoTA bias identification techniques on classical image datasets: Waterbirds [40] and CelebA [21], and further extend the evaluation by including a text dataset: CivilComments [5]. We also quantitatively confirm biases identified in ImageNet with respect to a foundation model.
  4. 4.
    We show that more robust methods can be obtained by devising constraints around the discovered biases. This includes subgroup constraints, training regularization, and zero-shot prompting.
Table 1: Comparison for Bias Identification approaches, based on fundamental differences in methods.
Method Root of bias identification Scoring bias candidates Fully auto Evaluation on large datasets Evaluation protocols employed
SpLiCE [4] disentangled representations Lasso solver Masking concepts from decomposition
B2T [15] mistakes in valid-set keywords frequency GroupDRO over bias-based groups
Lg [47] class-specificity score latent-space arithmetic GroupDRO over bias-based groups Train on bias-exploiting samples
ConceptDrift (ours) biases’ similarity to class weights latent-space arithmetic partially Validate on bias-exploiting samples (Sec. 5.1) GroupDRO over bias-based groups (Sec. 5.2) Bias-based training regularization (Sec. 5.3)
2 Related Work
Bias identification for generalization Machine learning methods naturally capture relevant factors needed to solve a task. Nevertheless, many times, models also capture shortcuts [12], that are helpful in solving the task, but are not fundamental or essential for it. These shortcuts represent spurious correlations or biases, that don’t always hold, and should not be used for reliable generalization outside the training distribution, often leading to degraded performance [29, 3, 13]. Prior works [15, 47, 4] have thus focused on identifying dataset biases, mainly through data analysis procedures. Kim et al. [15] identify concepts that are more prevalent in the misclassified examples in comparison with the correctly classified ones, while Zhao et al. [47] look for concepts that are specific to one class. In contrast, Bhalla et al. [4] identifies biases by training a linear layer on top of a disentangled representation of the data and having humans decide which features should not have been used for the task. For a meaningful comparison, we show in Tab. 1 existing bias identification methods, highlighting aspects we consider fundamental for the bias identification task.
Debiasing Debiasing and bias extraction techniques have become crucial in ensuring the fairness and accuracy of machine learning models [36], with extensive research dedicated to removing harmful biases across various domains. Some existing methods use bias annotations to train unbiased models, by means of group balanced subsampling [14], reweighting [33] or data augmentations [44]. In the absence of these annotations, other works [25, 20, 28] have proposed to first learn a biased model and then focus on its mistakes to train an unbiased one.
Fairness Fairness in has been extensively studied, with numerous approaches [8, 38] proposed to facilitate ethical research and to ensure equitable outcomes across different subpopulations. Most of those methods overlap with domain generalization and worst-group performance improvements. This is also a field where model interpretability plays a crucial role [32], as understanding how decisions are made can help in identifying and mitigating biases.
Invariant Learning Robustness to out-of-distribution changes can be obtained by enforcing that the learning model is invariant to different environments or domains [2, 16, 44]. But there are many cases where we don’t have access to such labeled environments and we must discover them. Approaches like [9] partition the data into subsets that maximally contradict an invariant constraint, and apply algorithms for distributional robustness, like groupDRO [33], on those subsets, called environments.
Model guidance with explanations Most similar to our bias-based regularization is the use of model explanations to guide learning [31, 35, 27, 42]. Using differentiable attribution methods like Grad-CAM [34], they constrain the attention of the trained models to only fall on the class instances, segmented at the input level. A limitation of these approaches is that they can’t prevent the model from learning biases that are not spatially separated from the class in the input (e.g. color, shape, or accessories of the subject).
@@ 3 Our Method
For a standard classification task, we aim to identify biases acquired by a model through training on a new dataset. By dataset biases we refer to class-independent concepts, whose presence in the samples greatly affects the class label distribution. By concepts we refer to words or expressions with a semantic load, that are most relevant in describing a given context. For instance, to extract concepts from an image dataset, we begin with the image captions, extract keywords from them, and use the keywords as concepts. Through biases of a classifier $f_{\theta}$ we refer to concepts causally unrelated with a class, whose presence in the input significantly changes the distribution of class probabilities predicted by $f_{\theta}$. We further classify them as positively correlated biases w.r.t. to a class $k$ if the presence of these biases in the input increases the probability of $f_{\theta}$ predicting the class $k$ and negatively correlated biases if they decrease it.
Setup We start with a dataset with samples $(x_{j},y_{j})$, in which we search for biases, through the representations offered by a foundation model $M$, capable of embedding both the input samples $x_{j}$ and the concepts $c_{i}$ into the $\mathbb{R}^{D}$ vector space. We further assume that their cosine similarity $cos(M(c_{i}),M(x_{j}))$ is higher when the concept $c_{i}$ is present in sample $x_{j}$ than when it is not. We first construct a set of concepts $(c_{i})_{1\leq i\leq q}$ in textual form, that are present in the training data. To spot the biases among them, we train a linear layer on top of the input data representations given by $M$. Next, through concept filtering, we identify concepts that are causally unrelated to the classes, revealing our bias candidates. Finally, we rank them based on how strongly they influence the predictions of this linear classifier and automatically decide the bias score cut-off. In the following equations, we use the dot product instead of the cosine similarity, as all the vectors we use are normalized. The main steps of our method (also revealed in Fig. 1, and in more detail in the Appx. C algorithm), are the following:
@@
Figure 2: The maximum distance between the reference line and the smoothed bias scores gives the threshold for our dynamic cut-off heuristic.
Step 1: Model training We train the weights using ERM [37] on our dataset of interest. Through learning, the class concepts (represented by the weights in the linear layer) naturally shift from the original initialization towards concepts that can effectively distinguish between class samples, given the current dataset, as visually presented in Fig. 1. This includes the classes themselves and their (positive) biases as well. At the end of the training, the weights can be expressed as a weighted sum of these concepts.
Step 2: Bias identification Using the learned weights, we continue to identify which concepts $c_{i}$ are either biases positively or negatively correlated with each class $k$. First, we filter out the class-related concepts. We then rank the remaining class-neutral concepts and select the top ones as biases based on a dynamic thresholding strategy, as follows:
a. Filter concepts We first aim to filter out concepts that are indeed related to a certain class, leaving only concepts that are causally unrelated to a class (we call them class-neutral concepts). These concepts are proper candidates for biases because they are not necessary for the recognition of a class (e.g. a forest background is well correlated with species of landbirds, but we want to prevent the model from using this causally unrelated correlation). The exact pipeline and methods are detailed in Sec. 4.
b. Rank class-neutral concepts For each class-neural concept $c_{i}$ and class $k$, we rank the concepts based on their similarity, using the following positive-bias score:
@@
Intuitively, we want concepts specific to one class but not to all the others. For this, for each class $k$, we select concepts highly correlated to it, while not correlated to at least one other class. For the negatively correlated biases, the score $s_{k,i}^{-}$ is similar, with the difference that we are taking into account the dissimilarity scores ($-w^{\top}_{k}M(c_{i})$).
Table 2: Validation procedure for three positively correlated biases found using ConceptDrift for CLIP ViT-L/14 fine-tuned on ImageNet. We evaluate the model’s capability to recognize a depicted (correct) class before and after the introduction of a biasing factor. For each prompt, $1000$ images are generated using FLUX.1-dev. We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the biasing factor is involved. We also observe a meaningful increase in the likelihood of having the induced class predicted, even though it is not illustrated.
Exploited Bias $\rightarrow$ Avg. Softmax Activations (%) Samples Predicted As (%)
Induced Class Correct Class Prompt Correct Class Induced Class Correct Class Induced Class
firemen (bias) $\rightarrow$ fire truck (class) peafowl a photo of a peafowl 99.9 0.0 100.0 0.0
firemen and a peafowl 10.6 (-89.3 ) 78.9 (+78.9) 5.3 (-94.7) 93.4 (+93.4)
reading a newspaper (bias) $\rightarrow$ crossword (class) Mexican hairless dog a photo of a mexican hairless dog 45.2 0.0 47.5 0.0
a man reading a newspaper in a chair with a mexican hairless dog in its lap 2.0 (-43.2) 21.6 (+21.6) 0.9 (-46.6) 36.6 (+36.6)
shrimp (bias) $\rightarrow$ American lobster (class) Bernese Mountain Dog a photo of a Bernese Mountain Dog 90.2 0.0 99.8 0.0
shrimp and pasta near a Bernese Mountain Dog 8.1 (-82.1) 14.8 (+14.6) 10.6 (-89.2) 37.2 (+37.2)
c. Dynamic thresholding towards biases We select the biases of each class using a dynamic threshold for the bias scores above. We keep the highest-ranked ones, both for positive and negative bias scores. We remove the null scores and then apply a mean filter with a window size of $r$ to smooth the curve of scores. We denote the scores obtained at this step with $(\overline{s}_{k,i})_{1\leq i\leq p}$, with $p=q-r+1$. We then select the top $m$ ones, as positive biases for class $k$, where the $m$ index is defined as:
@@
The intuition is that $m$ represents the index where the smoothed score curve deviates most from the straight reversed diagonal line, connecting $\overline{s}_{k,1}$ and $\overline{s}_{k,p}$ (see Fig 2).
Figure 3: Qualitative results depicting a series of biases uncovered employing our proposed ConceptDrift approach for CLIP ViT-L/14 fine-tuned on ImageNet [10]. All images are generated using FLUX.1-dev. In each scenario, a (correct) ImageNet class is depicted alongside a bias found to be positively correlated with a different (induced) class. Alongside each image, we provide the names of the correct and predicted classes, together with the softmax activations of their respective neurons. Though in most scenarios a single ImageNet class is clearly depicted, the model chooses to ignore it and predict a class that is not illustrated, clinging to the spurious correlations learned.

### v3

# WASP: A Weight-Space Approach to Detecting Learned Spuriousness

@@ Abstract
It is of crucial importance to train machine learning models such that they clearly understand what defines each class in a given task. Though there is a sum of works dedicated to identifying the spurious correlations featured by a dataset that may impact the model’s understanding of the classes, all current approaches rely solely on data or error analysis. That is, they cannot point out spurious correlations learned by the model that are not already pointed out by the counterexamples featured in the validation or training sets. We propose a method that transcends this limitation, switching the focus from analyzing a model’s predictions to analyzing the model’s weights, the mechanism behind the making of the decisions, which proves to be more insightful. Our proposed Weight-space Approach to detecting Spuriousness (WASP) relies on analyzing the weights of foundation models as they drift towards capturing various (spurious) correlations while being fine-tuned on a given dataset. We demonstrate that different from previous works, our method (i) can expose spurious correlations featured by a dataset even when they are not exposed by training or validation counterexamples, (ii) it works for multiple modalities such as image and text, and (iii) it can uncover previously untapped spurious correlations learned by ImageNet-1k classifiers. The code is available at https://github.com/bit-ml/WASP.
spurious correlation, ImageNet, waterbirds, CelebA, CivilComments, CLIP, foundation models, subpopulation shift
[…]
@@ 1 Introduction
Deep neural networks, and especially fine-tuned versions of foundation models, are commonly deployed in critical areas such as healthcare, finance, and criminal justice, where decisions based on spurious correlations (SCs) can have significant societal consequences (Angwin, Julia and Larson, Jeff and Mattu, Surya and Kirchner, Lauren, 2016; Caliskan et al., 2017). Even if the pretrained model has been validated by the community, the dataset leveraged in the fine-tuning process can, and usually does, imprint the model with new SCs. We present such an example in Fig. 1, where a CLIP model fine-tuned on ImageNet-1k labels the image as a ”Fire Truck” even in the complete absence of any fire truck or truck-related features and even given the presence of a peafowl (which is an ImageNet-1k class) right in front and center. This is due to the fact that, as our proposed WASP protocol discovered, the model spuriously correlates firefighters (which are not an ImageNet-1K class) with the ”Fire Truck” class.
Figure 1: Leveraging WASP’s identified spurious correlations, we generate synthetic images that fool strong pre-trained models.
Subpopulation shift setups (Wah et al., 2011b; Liu et al., 2015a; Borkan et al., 2019), featuring naturally occurring SCs, provide researchers with a controlled environment for studying SC detection and prevention. Within this context, recent efforts have begun to employ foundation models (Kim et al., 2024; Zhao et al., 2024; Bhalla et al., 2024) in their investigation protocols. Some of these efforts (Bhalla et al., 2024; Zhao et al., 2024) focused on finding SCs through data analysis, without referencing model predictions. While these findings are valuable, they may not always be relevant. For example, an SC may not impact a machine learning model if the correlation is harder to learn than the actual class itself. To this end, efforts such as that of (Kim et al., 2024) targeted learned spurious correlations, focusing on investigating validation samples that the model misclassifies. However, this error-based approach assumes that the validation set is in some sense exhaustive and that the validation data are already able to expose the SCs learned by the model, which may be true for subpopulation shift setups, but hardly guaranteed in any other general setup.
@@
Figure 2: Following WASP’s steps for the ImageNet-1k ”Fire Truck” class. In Step 1, during training, the classification weights $W$ drift from the initial class concept embedding $W^{0}$, outside the scope of relevant concepts, towards spuriously correlated ones. In Step 2, our method filters out class-related concepts and, using an embedding-space scoring system, ranks and automatically marks the highest-ranking class-neutral concepts as SCs.
@@
Different from prior work, we aim to overcome the limitations of data and error analysis, revealing spurious correlations learned by machine models even in scenarios in which the training and validation sets do not present counterexamples able to expose the model’s flaws. To this end we propose WASP, a Weight-space Approach to detecting learned Spurious correlations. Our overall framework is illustrated in Fig. 2 and further detailed in Sec. 3. We leverage the topology of foundation models such as CLIP and mGTE and observe that during the process of training, the weights of the final classification layer drift away from the textual representation of their associated class, towards identifying and prioritizing representations of spurious attributes. We propose a ranking system based on embedding-space arithmetic to extract concepts that factor in the activation of class neurons and delineate the highest-ranking concepts that lie outside the semantic scope of the classes, as spurious correlation.
Our work can be summarized as follows:
@@
    We introduce WASP, a weight-space approach to detecting learned spuriousness, offering a departure from the current error or data analysis methodology.
@@
    We show that our approach surpasses all state-of-the-art methods for identifying and naming spurious correlations in a given dataset. First, in its capacity to enhance the robustness of zero-shot models, and second, in its applicability to scenarios lacking spurious correlation counterexamples. Furthermore, in addition to evaluating our method on the established image datasets (Waterbirds, CelebA), we also validate its effectiveness on textual data (CivilComments).
@@
    We show that our method is able to expose previously untapped ImageNet-1k spurious correlations, and further proceed to show that multiple state-of-the-art models are affected by them, highlighting their inability to formulate robust definitions for the targeted classes.
2 Background
Machine learning methods naturally capture relevant factors needed to solve a task. However, models might also capture shortcuts (Geirhos et al., 2020), as correlations between non-essential features of the inputs and the label. These shortcuts represent spurious correlations, that don’t hold in a more general setup (e.g. using water features to classify waterbirds instead of focusing on the birds’ features), and should not be used for reliable generalization outside the training distribution, as they often lead to degraded performance (Quiñonero-Candela et al., 2008; Beery et al., 2018; Hendrycks et al., 2021).
SCs from error analysis
Approaches like B2T (Kim et al., 2024) rely exclusively on the validation samples, identifying which correlations between concepts are more prevalent in the misclassified examples. To catch SCs, B2T needs samples that oppose the strong correlations in the training set, thus leading to misclassification. To circumvent this limitation, DrML (Zhang et al., 2023b) manually builds a list of texts containing classes and concepts associations, that could potentially underline an SC. It forwards each such textual association through a classifier, learned on the image modality, and keeps as SCs the erroneously classified ones. In those approaches, the burden falls on the practitioners to come up with exhaustive samples or associations, failing to detect unexpected SCs. Differently, WASP focuses on analyzing the explicit learned weights of a model, covering all trainset samples. We thus extract the spuriously correlated concepts directly from the weights, bypassing the need for an exhaustive validation set or correlations candidates.
SCs from train data analysis
In SpLiCE (Bhalla et al., 2024) each image is decomposed into high-level textual concepts, searching next for concepts that are frequent for a certain class, but not for the others. LG (Zhao et al., 2024) relies on LLMs to propose concepts potentially correlated with each class, using image captions. Next, it uses CLIP (Radford et al., 2021) to estimate a class-specificity score for each concept, and highly scored concepts for a class w.r.t. the others are considered SCs. These methods focus on the concepts’ occurrences per class, making them prone to missing low-frequency concepts, as their presence can be drowned when averaging scores over a large dataset. Moreover, the SCs found through data analysis could be harder to learn than the class itself, so they are not necessarily imprinted in the model. In contrast, learned SCs (including error analysis revealed ones) must always be addressed, as they are, by definition, proven to impact the classifier. For this reason, WASP targets learned SCs by looking directly at the impact of the training set upon the model’s weights.
Manual interpretation of correlations
The method introduced in (Singla & Feizi, 2021) finds spurious features learned by a model, but it requires humans to manually annotate whether an image region is causal or not for a class. While this ensures a higher quality of the annotations, it also poses problems of scalability to large datasets. In contrast, WASP works fully automated, at scale, identifying SCs for each class in ImageNet-1k.
SCs from subpopulation shift setup Other previous works (Pezeshki et al., 2024; Liu et al., 2021a; Arefin et al., 2024; Zare & Nguyen, 2024) have focused on SC identification strictly within the context of subpopulation shifts. The particularity of this setup is that the training and validation sets always contain subsets of samples that oppose the strong spurious correlations of the dataset. With WASP, we do not have this limitation. Most of these methods (Pezeshki et al., 2024; Liu et al., 2021a; Zare & Nguyen, 2024) focus on first learning a strongly biased classifier and then either separate the samples of each class into two groups (Pezeshki et al., 2024; Zare & Nguyen, 2024) (one containing correctly classified examples and the other containing misclassified ones), or place higher weights on hard samples (Liu et al., 2021a), in order to balance the dataset. CoBalT (Arefin et al., 2024) on the other hand uses an unsupervised method for object recognition and then samples the dataset examples such that all object types are uniformly distributed in each class. Their result contains heatmaps overlays on images, which can offer insights to guide further manual SCs identification.
Fairness It is important to note that the proposed method can be utilized to evaluate the fairness of a given dataset and that we do conduct benchmarking on the CivilComments dataset, which encompasses racial and religious concerns. However, it is critical to emphasize that our approach is neither designed to measure nor address issues of fairness. Instead, our method is specifically developed to examine whether a given dataset imparts a clear definition of the featured classes to a model — namely, whether classifiers learn spurious correlations and confound class features with environmental features. Accordingly, our work is situated within the literature on subpopulation shift setups and we assess the quality of our proposed approach within this framework. Evaluating our approach on fairness benchmarks lies outside the scope of the current study, but may constitute a subject for subsequent research.
@@ 3 Our Method
For a standard classification task, we aim to identify spurious correlations learned by a model through training on a new dataset. By dataset’s SCs we refer to class-independent concepts, whose presence in the samples greatly affects the class label distribution. By concepts we refer to words or expressions with a semantic load, that are most relevant in describing a given context. Through SCs learned by a classifier $f_{\theta}$ we refer to concepts causally unrelated with a class, whose presence in the input significantly changes the distribution of class probabilities predicted by $f_{\theta}$. We further classify them as positively correlated concepts w.r.t. to a class $k$, if their presence in the input increases the probability of $f_{\theta}$ predicting the class $k$, and negatively correlated concepts if they decrease it.
Setup We start with a dataset with samples $(x_{j},y_{j})$, in which we search for SCs, through the representations offered by a foundation model $M$, capable of embedding both the input samples $x_{j}$ and the concepts $c_{i}$ into the $\mathbb{R}^{D}$ vector space. We further assume that their cosine similarity $cos(M(c_{i}),M(x_{j}))$ is higher when the concept $c_{i}$ is present in sample $x_{j}$ than when it is not. We first construct a set of concepts $(c_{i})_{1\leq i\leq q}$ in textual form, that are present in the training data. To spot the spuriously correlated ones among them, we train a linear layer on top of the input data representations given by $M$. Next, through concept filtering, we identify concepts that are causally unrelated to the classes, revealing our candidates. Finally, we rank them based on how strongly they influence the predictions of this linear classifier and automatically decide the SC score cut-off. In the following equations, we use the dot product instead of the cosine similarity, as all the vectors we use are normalized. The main steps of our method (also revealed in Fig. 2, and in more detail in the Appx. D algorithm), are the following:
@@
Figure 3: The maximum distance between the reference line and the smoothed scores gives the threshold for our cut-off heuristic.
Step 1: Model training We train the weights using ERM (Vapnik, 1999) on our dataset of interest. Through learning, the class concepts (represented by the weights in the linear layer) naturally shift from the original initialization towards concepts that can effectively distinguish between class samples, given the current dataset, as visually presented in Fig. 2. This includes not only the classes themselves but also concepts that are positively correlated with them. At the end of the training, the weights can be expressed as a weighted sum of classes and concepts.
Step 2: SCs identification Using the learned weights, we continue to identify which concepts $c_{i}$ are either positively or negatively correlated with each class $k$. First, we filter out the class-related concepts. We then rank the remaining class-neutral concepts and mark the top ones as SCs, based on a dynamic thresholding strategy, as follows:
a. Filter out class-related concepts Initially, we extract dataset’s concepts using existing tools. We first aim to filter out concepts that are indeed related to a certain class, leaving only concepts that are causally unrelated to a class (we call them class-neutral concepts). These concepts are proper candidates for SCs because they are not required for the recognition of a class (e.g. a forest background is well correlated with species of landbirds, but we want to prevent the model from using this causally unrelated correlation). The exact pipeline and methods are detailed in Sec. 4.
b. Rank class-neutral concepts For each class-neural concept $c_{i}$ and class $k$, we rank the concepts based on their similarity, using the following positive-SC score:
@@
Intuitively, we want concepts specific to one class but not to all the others. For this, for each class $k$, we select the concepts which starkly correlate to it, when compared to other classes. For the negatively correlated SCs, the score $s_{k,i}^{-}$ is similar, with the difference that we are taking into account the dissimilarity scores ($-w^{\top}_{k}M(c_{i})$).
Table 1: Validation procedure for three positively correlated concepts found using WASP for CLIP ViT-L/14 fine-tuned on ImageNet. We evaluate the model’s capability to recognize a depicted (correct) class before and after the introduction of an identified concept. For each prompt, $1000$ images are generated using FLUX.1-dev. We observe throughout all considered scenarios, a significant drop in the model’s capacity to identify the correct class when the selected concept is involved. We also observe a meaningful increase in the likelihood of having the induced class predicted, even though it is not illustrated.
Correct Exploited SC Prompt Avg. Softmax Activations (%) Samples Predicted As (%)
Class (Induced Class) Correct Class Induced Class Correct Class Induced Class
peafowl firemen $\bullet$ a photo of a peafowl 99.9 0.0 100.0 0.0
(fire truck) $\bullet$ firemen and a peafowl 10.6 (-89.3 ) 78.9 (+78.9) 5.3 (-94.7) 93.4 (+93.4)
Mexican reading a $\bullet$ a photo of a Mexican hairless dog 45.2 0.0 47.5 0.0
hairless newspaper $\bullet$ a man reading a newspaper in a chair 2.0 (-43.2) 21.6 (+21.6) 0.9 (-46.6) 36.6 (+36.6)
dog (crossword) with a Mexican hairless dog in his lap
Bernese shrimp $\bullet$ a photo of a Bernese Mountain Dog 90.2 0.0 99.8 0.0
Mountain (American $\bullet$ shrimp and pasta near a Bernese 8.1 (-82.1) 14.8 (+14.8) 10.6 (-89.2) 37.2 (+37.2)
Dog lobster) Mountain Dog
Table 2: Accuracy of various convolutional and transformer-based models trained on ImageNet-1k, on the data generated for Tab. 1. As with Fig. 1 and Fig. 4, we note that the performance of these models is significantly affected, even though the correct class is illustrated right in front and center while the predicted class is absent from the generated images. An exhaustive list is presented in Appx. B.
Prompt employed (correct class highlighted in bold, SC in red)
Model a photo of a peafowl firemen and a peafowl a photo of a Bernese Mountain Dog shrimp and pasta near a Bernese Mountain Dog
alexnet (Krizhevsky, 2014) $100.0$ $4.6$ (-95.4) $96.2$ $23.3$ (-72.9)
efficientnet_b1 (Tan & Le, 2019) $100.0$ $42.6$ (-57.4) $88.1$ $67.1$ (-21.0)
regnet_x_32gf (Radosavovic et al., 2020) $100.0$ $66.1$ (-33.9) $85.9$ $46.0$ (-39.9)
resnet50 (He et al., 2016) $100.0$ $30.1$ (-69.9) $73.9$ $54.5$ (-19.4)
resnext101_32x8d (Xie et al., 2017) $100.0$ $66.6$ (-33.4) $84.7$ $61.2$ (-23.5)
squeezenet1_1 (Iandola et al., 2016) $100.0$ $13.8$ (-86.2) $91.2$ $46.1$ (-45.1)
swin_b (Liu et al., 2021b) $100.0$ $81.5$ (-18.5) $95.2$ $72.6$ (-22.6)
vgg19_bn (Simonyan & Zisserman, 2015) $100.0$ $35.9$ (-64.1) $83.1$ $46.2$ (-36.9)
vit_l_16 (Dosovitskiy et al., 2021) $100.0$ $55.9$ (-44.1) $95.3$ $76.0$ (-19.3)
wide_resnet50_2 (Zagoruyko & Komodakis, 2016) $100.0$ $60.6$ (-39.4) $95.7$ $63.9$ (-31.8)
c. Dynamic thresholding We select the SCs of each class using a dynamic threshold for the scores above, keeping the highest-ranked ones. We remove the null scores and then apply a mean filter with a window size of $r$ to smooth the curve of scores. We denote the scores obtained at this step with $(\overline{s}_{k,i})_{1\leq i\leq p}$, with $p=q-r+1$. We then select the top $m$ ones, as positive SCs for class $k$, where the $m$ index is defined as:
@@
The intuition is that $m$ represents the index where the smoothed score curve deviates most from the straight reversed diagonal line, connecting $\overline{s}_{k,1}$ and $\overline{s}_{k,p}$ (see Fig. 3).
Figure 4: Qualitative results depicting a series of SCs, uncovered employing our proposed WASP approach for CLIP ViT-L/14 fine-tuned on ImageNet (Deng et al., 2009). All images are generated using FLUX.1-dev. In each scenario, a (correct) ImageNet class is depicted alongside a concept found to be positively correlated with a different (induced) class. We provide the names of the correct and predicted classes (which are specifically those induced by the chosen spuriously correlated concept). Though in most scenarios a single ImageNet-1k class is clearly depicted, the model ignores it and predicts a class that is not illustrated, clinging to the spurious correlations learned.